OpenStudy (anonymous):

2sin^2 b + 4cos(a+b)sinasinb + cos2(a+b) = ??

OpenStudy (amistre64):

any other information?

OpenStudy (anonymous):

options are sin2a , cos2a, sin2b, cos2b

OpenStudy (amistre64):

that last term cos^2 (a+b) or cos[2(a+b)] ?? the ascii is a bit vague since people tend to mistype exponent carrots

OpenStudy (anonymous):

cos[2(a+b)]

OpenStudy (amistre64):

seems like a nightmare of just applying identities .. cos(a+b) = cosa cosb - sina sinb if a=b the its cos^2 a - sin^2 a or 2 cos^2 a - sin a or cos a - 2sin^2 a

OpenStudy (anonymous):

so what will the answer be?

OpenStudy (amistre64):

dunno, have to run the gambit of simplifying the given with identities

OpenStudy (amistre64):

notice that sina sinb = cosa cosb - cos(a+b) but this might not be useful cos(a+b) can be replaced by cosa cosb - sina sinb which may reduce some things .... but at the moment its just a mess

OpenStudy (anonymous):

im trying it out, it just keeps getting more and more complicated.

OpenStudy (amistre64):

cos 2(a+b) = cos(2) cos(a+b) - sin(2) sin(a+b) which doesnt seem useful at the moment

OpenStudy (amistre64):

yeah its a pain :)

OpenStudy (anonymous):

i got it, i guess. we can take cos2(a+b) as cos(2a+2b) right?

OpenStudy (amistre64):

we can yes

OpenStudy (anonymous):

then, im getting cos2a

OpenStudy (amistre64):

chk with the wolf :)

OpenStudy (anonymous):

wolf? x)

OpenStudy (anonymous):

ohh. thanks. :D